The members of a club agreed to raise P150,000 for charity but 25 of them failed to pay their shares. As a result, each of the others had to pay P500 more. How many members has this club?
Answer by richwmiller(17219)

x*y=150000,
x=150000/y
(x-25)*(y+500)=150000
(150000/y-25)*(y+500)=150000
FOIL
150000/y*y+500*150000/y-25y-25*500=150000
500*150000/y-25y-25*500=0
20*150000/y-y-500=0
20*150000-y^2-500y=0
y^2+500y-3000000=0
(y-1500)*(y+2000)=0
y = 1500
x*y=150000
x = 100 members, y = 1500 original share